OpenStudy (fanduekisses):

How to evaluate log 27 81

OpenStudy (fanduekisses):

I set it to exponential form first right?

OpenStudy (anonymous):

\[\log_{81} 27\]

OpenStudy (fanduekisses):

I thought exp form was= \[27^{x}=81\]

OpenStudy (anonymous):

I was restating the problem, sorry for the confusion. You've got it right so far, do you know what to do next?

OpenStudy (fanduekisses):

Idk I'm guessing x= a fraction but idk I'm a little confused lol

OpenStudy (fanduekisses):

Oops btw I meant \[\log _{27}81\]

OpenStudy (anonymous):

Okay, that's a little different. That would be the log of 81 with base 27. You have the right exponential equation, though.

OpenStudy (fanduekisses):

Ok yeah sorry lol Thanks :)

OpenStudy (anonymous):

The next step should be to find the prime factorization of both 81 and 27. This way you'll know their common factors.

OpenStudy (fanduekisses):

9?

OpenStudy (anonymous):

That's a common factor, yes, but you need to know all the prime factors, and how many of them there are. Like for 27, the prime factorization would be 3 * 3 * 3

OpenStudy (fanduekisses):

Ohhh I see and for 81 3*3*3*3

OpenStudy (anonymous):

Exactly! And you can simplify these into 3^3 and 3^4, right? So now can you try to plug these into the exponential equation?

OpenStudy (fanduekisses):

How? like 1/3^3=1/3^4?

OpenStudy (anonymous):

Substituting 3^3 for 27 and 3^4 for 81. \[(3^3)^x=3^4\] Can you simplify that?

OpenStudy (fanduekisses):

the 3's cancel out? so 2x=4=2?

OpenStudy (anonymous):

I think you've got the right idea, except I'm not sure where you got the 2 from?

OpenStudy (fanduekisses):

ooooh I was thinking of it like 2x=4 divide 2 to both sides. But if that is wrong then what happens after the 3's cancel each other out?

OpenStudy (anonymous):

Oh, the first exponent is a 3, not 2.

OpenStudy (fanduekisses):

ohh

OpenStudy (fanduekisses):

so x= 4/3? :)

OpenStudy (anonymous):

Exactly, great job! :)

OpenStudy (fanduekisses):

Yay :D Thanks so much^_^ ♥

OpenStudy (anonymous):

No problem!